Bakerston Company is a manufacturing firm that uses job-order costing. The companys inventory balances were as follows at thRequired: a. Prepare a schedule of cost of goods manufactured in good form. (Do not round predetermined overhead rate. Inputb. Was the overhead underapplied or overapplied? By how much? (Do not round predetermined overhead rate. Input the amount asc. Prepare an income statement for the year. The company closes any underapplied or overapplied overhead to Cost of Goods Sol
Answer #1
Bakerston Company
Cost of goods Manufactured Report
For the Month of April
Beginning Raw Material Inventory $                                                11,600.00
Raw Material Purchases $                                            4,10,000.00
Less: Ending Raw Material $                                                15,900.00
Less: Indirect Material $                                                20,700.00
Direct Material Used=(A) $              3,85,000.00
Direct Labor=(B) $              3,39,000.00
Manufacturing Overhead Applied=(C )=(**) $              2,14,500.00
Total Current Manufacturing cost=(D )=(A)+(B)+(C ) $              9,38,500.00
Add: Beginning Work in Process=(E ) $                 32,600.00
Less: Ending Work in Process=(F) $                 14,300.00
Cost of goods Manufactured=(D)+(E )-(F) $              9,56,800.00
(**)=$261000*(14300/17400) 214500
Actual Overhead cost incurred
Indirect Material $                                                20,700.00
Indirect Labor $                                                74,000.00
Factory Utilities $                                                24,000.00
Factory Depreciation $                                            1,11,000.00
Total=(A) $                                            2,29,700.00
Overhead Applied=(B) $                                            2,14,500.00
Under Applied Overhead=(A)-(B) $                                                15,200.00
Bakerston Company
Income Statement
Sales Revenue=(A) $           12,84,000.00
Cost of goods sold
Beginning Finished goods inventory $                                            1,02,000.00
Add: Cost of goods manufactured $                                            9,56,800.00
Less: Ending Finished goods inventory $                                            1,25,000.00
Unadjusted cost of goods sold $              9,33,800.00
Less: Over applied Manufacturing Overhead $                 15,200.00
Total cost of good sold=(B) $              9,49,000.00
Gross Profit=(C )=(A)-(B) $              3,35,000.00
Less: Selling and Administrative Expenses=(D) $              2,80,000.00
Net Income from Operation=(C )-(D) $                 55,000.00
Selling & Administrative Expenses
Administrative Salaries $                                            1,51,000.00
Selling Cost $                                            1,11,000.00
Depreciation related to selling , general and administrative $                                                18,000.00
Total $                                            2,80,000.00